OpenStudy (anonymous):

simplify -x^2-2x+4=0 some technique other than square root formula

OpenStudy (anonymous):

use product of root(POR) and sum of root(SOR)

OpenStudy (anonymous):

would you do it for me, step by stepa

OpenStudy (anonymous):

POR=Product Of Root SOR=Sum Of Roots take -x^2-2x+4=0 change it to x^2+2x-4=0 (POR)A+B =2 (SOR)AB=-4 modify A+B=2 to A=2-B then insert to (2-B)B=-4 expand to 2B-B^2 =-4 then arrange to B^2-2B-4 then use formula for root got B=3.236 or B=-1.236 insert into A=2-B finally A=3.236 or A=-1.236

OpenStudy (anonymous):

POR=Product Of Root SOR=Sum Of Roots X^2-(SOR)X + (POR)=0 take -x^2-2x+4=0 change it to x^2+2x-4=0 same as x^2-(-2)-4=0 (POR)A+B =-2 (SOR)AB=-4 modify A+B=-2 to A=-2-B then insert to (-2-B)B=-4 expand to -2B-B^2 =-4 then arrange to B^2+2B-4 then use formula for root got B=-3.236 or B=1.236 insert into A=-2-B finally A=-3.236 or A=1.236

OpenStudy (anonymous):

the new one is the correct one

OpenStudy (anonymous):

POR=Product Of Root SOR=Sum Of Roots X^2-(SOR)X + (POR)=0 take -x^2-2x+4=0 change it to x^2+2x-4=0 same as x^2-(-2)x-4=0 (POR)A+B =-2 (SOR)AB=-4 modify A+B=-2 to A=-2-B then insert to (-2-B)B=-4 expand to -2B-B^2 =-4 then arrange to B^2+2B-4 then use formula for root got B=-3.236 or B=1.236 insert into A=-2-B finally A=-3.236 or A=1.236

OpenStudy (anonymous):

this one got no typing error~~ sorry for the mess

OpenStudy (anonymous):

thank you very much student 91 it helps a lot! jesusin